OpenStudy (anonymous):

D^4=a^4,where D=d/dx plzzzzz....find the four roots for D.(proceed this way D^4=(1)a^4 D=(1)^1/4 *a then use cis form in the place of 1.

OpenStudy (experimentx):

is a constant??

OpenStudy (anonymous):

yup

OpenStudy (experimentx):

D^3 = a^4t + k_1 D^2 = a^4t^2/2 + k_1t + k_2 ... ... integrate with t ...

OpenStudy (anonymous):

D^4=(cis 0)^1/4 *a

OpenStudy (experimentx):

do you have answer??

OpenStudy (anonymous):

yes...... a,-a,ia,-ia

OpenStudy (experimentx):

oh, \[ D^4 - a^4 = 0\] \[ (D^2 - a^2)(D^2 + a^2) = 0\]

OpenStudy (anonymous):

do u remember any formula like this.... cis(2k pi +1)theta /1.......something like this?

OpenStudy (experimentx):

Something like this?? \[ (D^2 + a^2)y = 0 => y = c_1 \cos(ax) + c_2 \sin(ax)\]

OpenStudy (anonymous):

if D^6=-1*a^6 ,then?

OpenStudy (experimentx):

is \[ D^2 = \frac{d^2y}{dx^2} \text { or } D^2 = (dy/dx)^2\]

OpenStudy (anonymous):

hmn.

OpenStudy (anonymous):

D^2=d^2/dx^2

OpenStudy (experimentx):

first one right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh na!!